2x 2 -(3+7i)x-(3-9i)=0 Find the value of x.Please also tell me what is the Euler form of complex No.

2x2-(3+7i)x-(3-9i)=0 Find the value of  x.Please also tell me what is the Euler form of complex No.

Grade:11

1 Answers

Vikas TU
4 years ago
2x2-(3+7i)x-(3-9i)=0
D = b^2 - 4ac 
b = -(3-7i) 
c = (9i-3)
D = (3-5i)^2
Now apply quadratic equation root formula 
x = 3i , (3+i)/2
Hope this helps 
Good luck